How do you find x if 4x+10=13x-8
Answer by rfer(16322)

4x+10=13x-8
our goal is to isolate x
subtract 10 from both sides
4x+10-10=13x-8-10
4x=13x-18
subtract 13x from both sides
4x-13x=-18
-9x=-18
we want x to be positive so mutiply both sides by "-"
-(-9x)=-(-18)
9x=18
divide both sides by 9
9x/9=18/9
x=2